OpenStudy (hihi):

What are the x-intercept(s) of the graph of y + 30 = x^2 + x? a. (−5, 0) and (6, 0) b. (−5, 0) and (−6, 0) c. (5, 0) and (−6, 0) d. (5, 0) and (6, 0)

OpenStudy (anonymous):

To find x intercept just put y=0, you get x^2+x-30=0, x^2+6x-5x-30=0, x(x+6)-5(x+6)=0 (x+6)(x-5)=0,x=-6or x=5

OpenStudy (hihi):

TY I see now.